OpenStudy (anonymous):

A certain metallic oxide contained 15% of moisture, 15% silica and the rest metallic oxide. after partial drying it contained 5% of moisture what is the % silica in the partially dried sample?

OpenStudy (anonymous):

@dpaInc @myininaya @mahmit2012@blues

myininaya (myininaya):

Hey @shameer1 ! I'm sorry but I don't know any chemistry .

OpenStudy (anonymous):

Drying will not eliminate metallic oxide, but it will certainly vaporize the water. With only 5% water remaining, I would say this is strictly a math problem: keep the silica/metallic oxide ratio the same before and after drying. Before: 15% water, 15% silica, 70% metallic ore After: 5% water, x % silica, y % metallic ore System of equations: x + y = 95 4.67x = y I used substitution, and got x ~ 16.8%, so y ~ 78.2% If anyone else knows better, please post.

OpenStudy (anonymous):

initial percentage of moisture=15% " " " silica =15% " " " metallic oxide=70% %ratio of silica to metallic oxide=15/70=3/14 AFTER PARTIAL DRYING % of moisture = 5% let percentage of silica be x percentage of metallic oxide = 95-x then, x / 95-x =3/14 => 14x= 285-3x =>17x=285 x=16.67% approximately THE ANSWER IS ~16.67%